If E and F are two disjoint events in S with P(E) = 0.6 and P(F) = 0.1, find P(E ∪ F), P(EC), P(E ∩ F), P((E ∪ F)C), and P((E ∩ F)C).

P(E ∪ F) =
P(EC) =
P(E ∩ F) =
P((E ∪ F)C) =
P((E ∩ F)C) =

Answer by ikleyn(53763)

.
If E and F are two disjoint events in S with P(E) = 0.6 and P(F) = 0.1, find P(E ∪ F), P(EC), P(E ∩ F), P((E ∪ F)C), and P((E ∩ F)C).

P(E ∪ F) = P(E) + P(F) = 0.6 + 0.1 = 0.7.
	 	 
P(EC)          = 1 - P(E) = 1 - 0.6 = 0.4.

P(E ∩ F) = 0.

P((E ∪ F)C) = 1 - P(E) - P(F) = 1 - 0.6 - 0.1 = 0.3.

P((E ∩ F)C) = 1 - P(E ∩ F) = 1 - 0 = 1.